OpenStudy (anonymous):

Complete the following identities 1. Sin^2Q+.......=1 2. sinQ~=...... cosQ

OpenStudy (ikram002p):

ok so first u need to kno this identity \(\sin^2 \theta +cos^2 \theta =1\)

OpenStudy (ikram002p):

how ever what is number 2 u have ?